# Critical Evaluation of Studies Alleging Evidence for Technosignatures in the POSS1-E Photographic Plates

### v1

@@ V Frequency of occurrence in Earth’s shadow
Another important problem with the in-shadow frequency calculation is that set $V$ was derived from $S$ which, as we have seen, was determined by Solano et al. (2022) to consist primarily of objects that have not been distinguished from catalog stars and scan defects. At least 95% (i.e., $1-N_{\rm R}/N_{\rm V}$) of SPFs in $V$ were discarded by Solano et al. (2022) to produce the aggressively-vetted subset $R$ (see Table 1). And yet, $V$ was used instead of $R$ to compute the in-shadow frequency of SPFs.
@@ VI Temporal distribution and alleged correlation with nuclear tests
Bruehl and Villarroel (2025) sought to relate the timing of SPFs in $V$ and the timing of nuclear weapons tests. They defined the time window of their study as spanning November 19, 1949 to April 28, 1957, inclusive. Of the 2,718 days in this period, “observed transients” (SPFs) were found on only 310 days according to Bruehl and Villarroel (2025). However, an examination of the POSS1-E plates set reveals that the 937 plate observations were conducted on a total of only 380 days, of which 368 days (906 plates) fall within this study time period. The analysis that follows brings the number of observation days in dataset $V$ down to a maximum of 312. Based on $W$ and $R$, it is also likely that nearly all of the POSS1-E plates sampled to construct $V$ contained SPFs, suggesting that the reported correlation between SPFs and nuclear tests is almost entirely determined by the Palomar telescope observation schedule. We have plotted the dates of the nuclear tests, as well as the POSS1-E plate observation days (“plate days”), in Figure 10.
@@ VI.2 Normalization by the true observation window
SPFs are reported in the $V$ dataset (Villarroel et al., 2025a) on 310 days out of the 2,718 days in their study period. Bruehl and Villarroel (2025) find that 255 of these 310 days occur outside any nuclear testing window and 54 days occur inside a nuclear testing window; the testing window is defined by Bruehl and Villarroel (2025) as a day of nuclear testing $\pm 1$ day. In order to test for significant correlation between the detection of SPFs and the nuclear testing windows, Bruehl and Villarroel (2025) compared these numbers (255 days and 54 days, respectively) to the days with and without detected SPFs, that occur within the full time period of 2,718 days. In reality, the POSS1 survey did not record observations every single day out of this time period (Minkowski and Abell, 1963; STScI, 2026). There are plates available for only 380 days, of which 368 fall within the study period.
In addition, we argue that the conservative normalization of 368 days counted within the study period during which plates are available, can be further reduced if we account for the lack of SPF sampling in the southern hemisphere. The footprints of 31% of the 937 POSS1-E plates reside entirely in the southern celestial hemisphere without overlapping the celestial equator. But as explained in Section III, the datasets $R$, $W$ and $V$ contain SPFs overwhelmingly located in the northern celestial hemisphere. Set $W$ contains less than 0.1% of SPFs in the southern celestial hemisphere, while $R$ contains 1.4% and $V$ also contains 1.4% (see Section III and Appendix A). The southern hemisphere SPFs in these datasets reside overwhelmingly in equatorial plates that were acquired on the same day as at least one northern hemisphere plate, and so they do not contribute to the count of total unique observation days. This discrepancy between the northern versus southern hemisphere distribution of SPFs suggests that the plates whose footprints fully reside in the southern hemisphere were not sampled in search of SPFs for set $A$ and its subsets ($S$, $P$, $W$, $R$, and $V$) (see Section III and Appendix A), and the corresponding southern hemisphere-only observation days should be removed from the total count of observation days (Figure 10).
@@ VI.3 Correlation between POSS1 observation schedule and nuclear test schedule
Table 3: Left table: reproduction of Table 1 from Bruehl and Villarroel (2025). Middle and right tables: our day counts obtained after properly normalizing by the observation window. We have used the most vetted dataset $R$ instead of $V$. Due to smaller statistics overall, $R$ contains only 290 days where SPFs were found as opposed to 310 days found by Bruehl and Villarroel (2025) in the $V$ dataset. The nuclear testing window is defined in Bruehl and Villarroel (2025) as a day of nuclear testing $\pm 1$ day. Percentages in parenthesis show the conditional probabilities of observing (or not observing) an SPF given the presence (or absence) of a nuclear test on that day. Observation days relevant to this study are the days on which plates are sampled (“plate days” in Figure 10) solely from the northern celestial hemisphere and/or overlapping the celestial equator.
@@
Overall, it is not surprising to find a correlation between days on which SPFs were recorded and nuclear test windows considering that (i) SPFs are reported by Bruehl and Villarroel (2025) on at least 84% (probably 99%) of the days on which the POSS1 survey was observing; and (ii) the POSS1 observation schedule has a non-negligible (35%) likelihood of falling within nuclear test windows compared to outside. Moreover, normalizing by the effective POSS1 observation days instead of the full time period significantly dilutes the significance of the statistical effect reported between SPF detection and nuclear testing. We speculate that the correlation may be due to seasonality, since both the timing of astronomical observations and nuclear testing rely on clear weather conditions, but further investigation would be required to fully explain it.
@@ VII Discussion
Research related to SETI and UAP is commonly concerned with searching for unusual signals against a background that is contaminated with artifacts and other sources of noise. For these fields to advance, it is essential that analyses and conclusions are preceded by rigorous data validation. This step is essential even in cases where it is extremely difficult. Solano et al. (2022) constructed the dataset $R$ ($N=5,399$) after removing objects that could not be readily distinguished from astronomical objects, plate artifacts, and scan artifacts. A logical next step is to determine whether any of these features represent an optical transient: that is, whether any originated from a light source in front of the telescope, with a corresponding optical path. The lack of data validation in Villarroel et al. (2025a) and Bruehl and Villarroel (2025) has led to the erroneous assumptions and reasoning, contradictions, and mistaken conclusions that we have described in this paper.
@@ VII.1 Analyzing unvalidated data
The most important problem with the Villarroel et al. (2025a) and Bruehl and Villarroel (2025) studies is that the dataset $V$ and its subsets have not been validated in ways that are necessary for their use in a scientific study. In reference to the dataset $V$ of “carefully selected transient samples,” the authors concede that “this sample has not been visually inspected. As such, it is expected to contain a substantial number of false positives, including clustered artifacts such as edge fingerprints or other plate defects that contaminate our sample.” The basic question of “what do the data points represent?” has not been answered. We have argued that this is absolutely required for anything of value to be extracted from these data.
We suggest that it is probably essential to perform detailed experiments to record optical flashes with a similar or identical instrument and optical medium to gain confidence that optical flash candidates can be distinguished from plate defects. Based on previous work (see Section II) this requires the identification of a robust set of criteria for microscopic evaluation of individual features in the emulsion of glass plates. A parallel task is to ensure that each optical flash candidate is not related to any known fast astronomical transient or other known sources of optical flashes (see for example Hudec (1993)). This work may benefit from a microscopic examination of known high-altitude satellite glints in photographic plates post-dating POSS1. In addition, it would be important to study the processes that tend to degrade and contaminate photographic plates, resulting in additional plate artifacts. As previously noted, a morphological analysis with a more stringent symmetry filter could be used to remove additional artifacts, in order to highlight a smaller subset of $R$ that could serve as the focus of these investigations.
@@ VII.2 Importance of characterizing and addressing background variations
Before performing tests of statistical significance to understand the relationship to terrestrial shadowing, the null probability of alignments, and the correlation with nuclear tests, it is vital to examine and understand what the dataset contains: what its data points represent. If, as in the present case, the composition of the dataset is not known, the main sources of variation can at least be explored. A further essential step is to characterize and model these sources of variation (e.g., the overdensity in regions near the Galactic plane) and account for them before examining the relationship to other factors. In the present study, we have shown that the primary processes generating SPFs are very likely not optical transients; these processes should be modeled and addressed before assessing second-order effects and searching for rare events like optical transients.
@@ VII.3 Expected spatial distribution of plate artifacts
This is prominently revealed by the attention paid to statistical tests of significance, to illustrate a nonrandom distribution, such as a deficit of SPFs in Earth’s shadow and SPF alignments. Since we do not understand in detail the processes that can generate plate artifacts, it is not clear that a random distribution is the correct null hypothesis. A templating process (see our Section IV.3) could produce plate artifacts exhibiting alignments; for example, the shape created by SPFs in part A of Figure 4 has sharply-defined edges made from dense clusters of SPFs. A relative deficit of SPFs in a tiny fraction of the sky (e.g., the simulated terrestrial shadow at GSO covered just 1.5% of the sampled POSS1-E plates) could be entirely explained by excesses of SPFs that may have occurred in plates strongly affected by degradation (see Section IV.1), or conversely by deficits arising from features like the large-scale voids visible in Figure 8 (see Section IV.4).
@@ VII.4 Inconsistent statements concerning data preparation
Villarroel et al. (2025a) contains statements about the extent to which $V$ and its subsets were filtered to remove spurious detections and catalog objects, which are often inconsistent with each other or inconsistent with descriptions in Solano et al. (2022) about the preparation of the parent sets $A$ and $S$. In Section 2 of Villarroel et al. (2025a), the authors write “we shall use carefully selected transient samples in Solano et al. (2022), which… have been matched to several modern surveys to remove variable stars, asteroids, and comets.” But $V$ and its subsets are nowhere defined in Solano et al. (2022), which states that asteroids and variable stars were removed as part of the construction of $R$ ($N=5,399$), only after removing star-proximate SPFs from $S$ (see Table 1); this leaves a remainder of $N=9,171$, which is clearly too small to be the source of $V$. The removal of “comets” is mentioned nowhere in Solano et al. (2022). These and other inconsistent statements have made it impossible to accurately reconstruct the definitions of datasets. We have highlighted additional inconsistent remarks about the dataset $V$ and its derivatives in Appendix A.
VII.5 Circular reasoning and tautological claims
Circular or tautological reasoning involves assuming the conclusion of an argument as part of its premises. Villarroel et al. (2025a) and Bruehl and Villarroel (2025) make widespread use of the word “transient” to describe what we have termed “SPFs” (Selected POSS1 Features). The word “transient” implies that the features in question are in fact images of short-lived light sources, before this has been demonstrated. For example, consider this statement from page 2 of Villarroel et al. (2025a):
@@ VII.5 Circular reasoning and tautological claims
As we have discussed in Section VI, the correlation between SPFs and nuclear tests appears to be determined largely or entirely by the Palomar telescope observation schedule. This correlation may therefore be related to seasonal changes and other conditions that are favorable to both activities. Section V also showed that set $R$, the most aggressively vetted set of SPFs, has no deficit of SPFs in Earth’s shadow.
@@ VIII Conclusions
    We conducted a review of optical transient searches in photographic plates, in the context of historical searches for the optical counterparts of Gamma Ray Bursts (GRBs). Following two decades of work, researchers were not able to make a confident identification; even tentative identification of optical flashes required close microscopic analysis, which Villarroel et al. (2025a) did not perform. This and preceding work (Villarroel et al., 2021, 2022; Solano et al., 2022, 2024) have not heeded the primary lessons from this literature. If robust criteria for identifying optical flashes in archival plates are eventually identified, this leaves the problem of how to rule out alternative sources of optical flashes such as astronomical fast transients and flashing lights on aircraft.
@@
    The on-sky distribution of SPFs in $W$ reveals a window of right ascension in which there is a marked deficit of SPFs not present in the POSS1-derived $M$ dataset of likely celestial objects. We have also highlighted a void above the Galactic plane in the western celestial hemisphere in set $R$. We have pointed out the importance of characterizing and modeling sources of spatial variation at all scales in order to identify sources of variation influencing the in-shadow frequency and which are unrelated to the shadow itself. We find no significant in-shadow deficit of SPFs from the aggressively-filtered set $R$.
@@
    After normalizing properly by the number of observation days (the days on which plates were exposed, rather than the total span of time in the study window), the correlation reported in Bruehl and Villarroel (2025) between SPFs and nuclear tests can be explained by the incidental correlation between nuclear tests and the POSS1-E observation schedule. This correlation may, in turn, depend on factors that influence both of these activities, such as seasonal effects.
@@
    We have also highlighted examples of tautological claims and circular reasoning. Of special note is an argument invoking the consequences of explanatory hypotheses (i.e., an expected deficit of SPFs in the Earth’s shadow) to provide justification for confidence in the validity of measurements (i.e., that the SPFs represent images of real objects in space).
In summary, our study suggests that before proceeding further with the research program described in Villarroel et al. (2025a) and Bruehl and Villarroel (2025), far more work is needed to clearly define, understand, and validate the datasets under scrutiny. This work has been helpful, however, as motivating the contemporary search for technosignatures near Earth using specially-designed instrumentation, such as telescope arrays for triangulating the positions and characterizing the kinematics of anomalous interplanetary objects Villarroel (2024); Villarroel et al. (2025c). Although in some ways more complicated owing to the abundance of human spacecraft in orbit around Earth, this direction can avoid some of the challenges and pitfalls plaguing searches in historical photographic plates, as exemplified by GRB research after 1996.

### v2

@@ V Frequency of occurrence in Earth’s shadow
Another important problem with the in-shadow frequency calculation is that set $V$ was derived from $P$ which, as we have seen, was determined by Solano et al. (2022) to consist primarily of objects that have not been distinguished from catalog stars and scan defects. At least 95% (i.e., $1-N_{\rm R}/N_{\rm V}$) of SPFs in $V$ were discarded by Solano et al. (2022) to produce the aggressively-vetted subset $R$ (see Table 1). And yet, $V$ was used instead of $R$ to compute the in-shadow frequency of SPFs.
@@ VI Temporal distribution and alleged correlation with nuclear tests
Bruehl and Villarroel (2025) sought to relate the timing of SPFs in $V$ and the timing of nuclear weapons tests. The time window of the study is defined as spanning November 19, 1949 to April 28, 1957, inclusive. Of the 2,718 days in this period, “observed transients” (SPFs) were found on only 310 days according to Bruehl and Villarroel (2025). However, an examination of the POSS1-E plates set reveals that the 937 plate observations were conducted on a total of only 380 days, of which 368 days (906 plates) fall within this study time period. The analysis that follows brings the number of observation days in dataset $V$ down to a maximum of 312. We have plotted the dates of the nuclear tests, as well as the POSS1-E plate observation days (“plate days”), in Figure 10.
@@ VI.2 Normalization by the true observation window
SPFs were reported on 310 days out of the 2,718 days in the study period (Bruehl and Villarroel, 2025), of which 255 days occur outside any nuclear testing window and 54 days occur inside a nuclear testing window. The testing window is defined by Bruehl and Villarroel (2025) as a day of nuclear testing $\pm 1$ day. In order to test for significant correlation between the detection of SPFs and the nuclear testing windows, Bruehl and Villarroel (2025) compared these numbers (255 days and 54 days, respectively) to the days with and without detected SPFs, that occur within the full time period of 2,718 days. In reality, the POSS1 survey did not record observations every single day out of this time period (Minkowski and Abell, 1963; STScI, 2026). There are plates available for only 380 days, of which 368 fall within the study period.
In addition, we argue that the conservative normalization of 368 days counted within the study period during which plates are available, can be further reduced if we account for the lack of SPF sampling in the southern hemisphere. The footprints of 31% of the 937 POSS1-E plates reside entirely in the southern celestial hemisphere without overlapping the celestial equator. But as explained in Section III, the datasets $R$, $W$, and $V$ contain SPFs overwhelmingly located in the northern celestial hemisphere. Set $W$ contains less than 0.1% of SPFs in the southern celestial hemisphere, while $R$ contains 1.4% and $V$ also contains 1.4% (see Section III and Appendix A). The southern hemisphere SPFs in these datasets reside overwhelmingly in equatorial plates that were acquired on the same day as at least one northern hemisphere plate, and so they do not contribute to the count of total unique observation days. This discrepancy between the northern versus southern hemisphere distribution of SPFs suggests that the plates whose footprints fully reside in the southern hemisphere were not sampled in search of SPFs for set $A$ and its subsets ($S$, $P$, $W$, $R$, and $V$) (see Section III and Appendix A), and the corresponding southern hemisphere-only observation days should be removed from the total count of observation days (Figure 10).
@@ VI.3 Correlation between POSS1 observation schedule and nuclear test schedule
Table 3: Left table: reproduction of Table 1 from Bruehl and Villarroel (2025). Middle and right tables: our day counts obtained after properly normalizing by the observation window. We have used the most vetted dataset $R$ instead of $V$. Due to smaller statistics overall, $R$ contains only 289 days where SPFs were found as opposed to 310 days found by Bruehl and Villarroel (2025) in the $V$ dataset. The nuclear testing window is defined in Bruehl and Villarroel (2025) as a day of nuclear testing $\pm 1$ day. Percentages in parenthesis show the conditional probabilities of observing (or not observing) an SPF given the presence (or absence) of a nuclear test on that day. Observation days (“plate days” in Figure 10) relevant to this study are the days on which at least one plate is sampled from the northern celestial hemisphere and/or overlapping the celestial equator.
@@
Overall, it is not surprising to find a correlation between days on which SPFs were recorded and nuclear test windows considering that (i) SPFs are reported by Bruehl and Villarroel (2025) on at least 84% (probably 99%) of the days on which the POSS1 survey was observing; and (ii) the POSS1 observation schedule has a non-negligible (35%) likelihood of falling within nuclear test windows compared to outside. Moreover, normalizing by the effective POSS1 observation days instead of the full time period significantly dilutes the significance of the statistical effect reported between SPF detection and nuclear testing. We speculate that the correlation may be due to seasonality, since both the timing of astronomical observations and nuclear testing rely on clear skies and low-wind weather conditions, but further investigation will be required to explain it.
@@ VII Discussion
Research related to SETI and UAP is commonly concerned with searching for unusual signals against a background that is contaminated with artifacts and other sources of noise. For these fields to advance, it is essential that analyses and conclusions are preceded by rigorous data validation. This step is essential even in cases where it is extremely difficult. Solano et al. (2022) constructed the dataset $R$ ($N=5,399$) after removing objects that could not be readily distinguished from astronomical objects, plate artifacts, and scan artifacts. A logical next step is to systematically examine this dataset to determine whether any of these features are due to emulsion or other defects and, finally, to determine if any that remain represent true optical sources: that is, whether any originated from a light source in front of the telescope, with a corresponding optical path. The lack of data validation in Villarroel et al. (2025a) and Bruehl and Villarroel (2025) has led to some of the erroneous assumptions and reasoning, contradictions, and mistaken conclusions that we have described in this paper.
@@ VII.1 Analyzing unvalidated data
The underlying problem with the Villarroel et al. (2025a) and Bruehl and Villarroel (2025) studies is that the dataset $V$ and its subsets have not been validated in ways that are necessary for their use in a scientific study. In reference to the dataset $V$ of “carefully selected transient samples,” it is conceded that “this sample has not been visually inspected. As such, it is expected to contain a substantial number of false positives, including clustered artifacts such as edge fingerprints or other plate defects that contaminate our sample” (Villarroel et al., 2025a). The basic question of “what do the data points represent?” has not been answered. We have argued that this is absolutely required for any robust scientific result to be derived from these data.
We suggest that it is probably essential to perform detailed experiments to record optical flashes with a similar or identical instrument and optical medium to gain confidence that optical flash candidates can be distinguished from plate defects. Based on previous work (see Section II) this requires the identification of a robust set of criteria for microscopic evaluation of individual features in the emulsion of glass plates. A parallel task is to ensure that each optical flash candidate is not related to any known fast astronomical transient or other known sources of optical flashes (see for example Hudec (1993)). This body of work may benefit from a microscopic examination of known high-altitude satellite glints in photographic plates post-dating POSS1. In addition, it would be important to study the processes that tend to degrade and contaminate photographic plates, resulting in additional plate artifacts. As previously noted, a morphological analysis with a more stringent symmetry filter could be used to remove additional artifacts, in order to highlight a smaller subset of $R$ that could serve as the focus of these investigations.
@@ VII.2 Importance of characterizing and addressing background variations
Before performing tests of statistical significance to understand the relationship to terrestrial shadowing, the null probability of alignments, and the correlation with nuclear tests, it is vital to examine and understand what its data points represent. If, as in the present case, the composition of the dataset is not known, the main sources of variation can at least be explored. A further essential step is to characterize and model these sources of variation (e.g., the overdensity in regions near the Galactic plane, the gaps along bands of right ascension) and account for them before examining the relationship to other factors. In the present study, we have shown that the primary processes generating SPFs are very likely not optical transients; these processes should be modeled and addressed before assessing second-order effects and searching for rare events like optical transients.
@@ VII.3 Expected spatial distribution of plate artifacts
This is prominently revealed by the attention paid to statistical tests of significance, to illustrate a nonrandom distribution, such as a deficit of SPFs in Earth’s shadow and SPF alignments. Since we do not understand in detail the processes that can generate plate artifacts, it is not clear that a random distribution is the correct null hypothesis. A templating process (see our Section IV.3) could produce plate artifacts exhibiting alignments; for example, the shape created by SPFs in part A of Figure 4 has sharply-defined edges made from dense clusters of SPFs. A relative deficit of SPFs in a tiny fraction of the sky (e.g., the simulated terrestrial shadow at GSO covered just 1.5% of the sampled POSS1-E plates) could be entirely explained by excesses of SPFs that may have occurred in plates strongly affected by degradation, or conversely by deficits arising from features like the large-scale voids visible in Figure 8 (see Section IV.4).
@@ VII.4 Inconsistent statements concerning data preparation
Villarroel et al. (2025a) contains statements about the extent to which $V$ and its subsets were filtered to remove spurious detections and catalog objects, which are often inconsistent with each other or inconsistent with descriptions in Solano et al. (2022) about the preparation of the parent sets $A$ and $S$. In Villarroel et al. (2025a), Section 2 specifies that it “shall use carefully selected transient samples in Solano et al. (2022), which… have been matched to several modern surveys to remove variable stars, asteroids, and comets.” But $V$ and its subsets are nowhere defined in Solano et al. (2022), which states that asteroids and variable stars were removed as part of the construction of $R$ ($N=5,399$), only after removing star-proximate SPFs from $S$ (see Table 1); this leaves a remainder of $N=9,171$, which is clearly too small to be the source of $V$. The removal of comets is mentioned nowhere in Solano et al. (2022). These and other inconsistent statements have made it impossible to accurately reconstruct the definitions of datasets. We have highlighted additional inconsistent remarks about the dataset $V$ and its derivatives in Appendix A.
VII.5 Circular reasoning
Circular reasoning involves assuming the conclusion of an argument as part of its premises. Villarroel et al. (2025a) and Bruehl and Villarroel (2025) make widespread use of the word “transient” to describe what we have termed “SPFs” (Selected POSS1 Features). The word “transient” implies that the features in question are in fact images of short-lived light sources, before this has been demonstrated. For example, consider this statement from page 2 of Villarroel et al. (2025a):
@@ VII.5 Circular reasoning
The first example suggests that if there is a deficit in the shadow, the data should be understood as representing valid measurements (images of objects in front of the telescope) and that these measurements represent glinting objects in orbit around Earth; and if there isn’t, then the data is taken to represent neither of these.
The second example suggests that if there is correlation with nuclear testing, then the data should be understood as representing valid measurements, and that these measurements represent UAP; likewise, the absence of a correlation would be taken to imply neither of these.
As discussed above, Section V showed that set $R$, the most aggressively vetted set of SPFs, has no deficit of SPFs in Earth’s shadow. In Section VI, the correlation between SPFs and nuclear tests is revealed to be determined largely or entirely by the Palomar telescope observation schedule.
To escape the circular reasoning, the underlying validity of the measurements must be established independently of the inferential analysis, which is in turn cited as support for the conclusion.
@@ VIII Conclusions
    We conducted a review of optical transient searches in photographic plates, in the context of historical searches for the optical counterparts of Gamma Ray Bursts (GRBs). Following two decades of work, even tentative identification of optical flashes required close microscopic analysis of glass plates, which Villarroel et al. (2025a) did not perform. This and preceding work (Villarroel et al., 2021, 2022; Solano et al., 2022, 2024) have not heeded the primary lessons from this literature. If robust criteria for identifying optical flashes in archival plates are eventually identified, this leaves the problem of how to rule out alternative sources of optical flashes such as astronomical fast transients and flashing lights on aircraft.
@@
    The on-sky distribution of SPFs in $W$ reveals windows in right ascension in which there is a marked deficit of SPFs not present in the POSS1-derived $M$ dataset of likely celestial objects. We have also highlighted a void above the Galactic plane in the western celestial hemisphere in set $R$. We have pointed out the importance of characterizing and modeling sources of spatial variation at all scales in order to identify sources of variation influencing the frequency in Earth’s shadow and which are unrelated to the shadow itself. We find no significant in-shadow deficit of SPFs from the aggressively-filtered set $R$.
@@
    After normalizing properly by the number of observation days (the days on which plates were exposed, rather than the total span of time in the study window), we show that SPFs are found on 310 of the 312 observations days, and thus the correlation reported in Bruehl and Villarroel (2025) between SPFs and nuclear tests can be explained by the incidental correlation between nuclear tests and the Palomar telescope observation schedule during the POSS1 survey. This correlation may, in turn, depend on factors that influence both of these activities, such as seasonal effects.
@@
    We have also highlighted examples of circular reasoning. Of special note is an argument invoking the consequences of explanatory hypotheses (i.e., an expected deficit of SPFs in the Earth’s shadow) to provide justification for confidence in the validity of measurements (i.e., that the SPFs are images of real objects in front of the telescope) as well as the conclusion (i.e., that these images represent glinting objects in outer space).
In summary, our study suggests that before proceeding further with the research program described in Villarroel et al. (2025a) and Bruehl and Villarroel (2025), far more work is needed to clearly define, understand, and validate the datasets under scrutiny. However, this research program is pursuing an intriguing question. The effort so far has been instrumental in motivating the contemporary search for technosignatures near Earth using specially-designed instrumentation, such as telescope arrays for triangulating the positions and characterizing the kinematics of anomalous interplanetary objects (Villarroel, 2024; Villarroel et al., 2025c). Although in some ways more complicated owing to the abundance of human spacecraft in orbit around Earth, this direction will avoid many of the challenges and pitfalls plaguing searches in historical photographic plates, and potentially enable the kind of rapid progress made by GRB optical counterpart research after 1996.